Sarah has 16 pounds of soup that is 1/4 chicken broth. How much chicken broth must she add to increase the chicken broth content to 1/3 ?
Answer by VFBundy(438)

Broth/soup = 1/4. We are given the soup is 16 lbs., so to get 1/4 broth, that means the broth must be 4 lbs. (Because broth/soup = 4/16 = 1/4.)

To find out how much broth we must add to get a soup that is 1/3 broth:

Broth/soup = (4+x)/(16+x) = 1/3

Cross-multiply:
3(4+x) = 1(16+x)

12 + 3x = 16 + x

2x = 4

x = 2

Sarah must add 2 lbs. of chicken broth. This will give her 6 total lbs. of broth and 18 total lbs. of soup, for a soup that is 1/3 broth.